OpenStudy (anonymous):

Prove: 24|(5^(2n)-1) for every positive integer n.

OpenStudy (anonymous):

And it needs to be proved using induction!

ganeshie8 (ganeshie8):

where are u stuck

OpenStudy (anonymous):

I showed for n=1 this way: 5^(2*1)-1= 24m 25-1=24m 24=24m m=1 Since m is an integer, it is true for n =1. Then I assume it is true for all n = k. I am stuck on the inductive hypothesis part where i need to prove for n = k+1

ganeshie8 (ganeshie8):

assume the statement is true for \(n=k\) : \(\large 5^{2k} - 1 = 24q\) you need to provethe statement is true for \(n=k+1\)

ganeshie8 (ganeshie8):

for \(n = k+1\) : \(5^{2(k+1)} - 1 = 5^2(5^{2k} - 1) + 24 = 5^2(24q) + 24 = 24(5^2q+1) = 24r\) QED.

OpenStudy (anonymous):

Yes i proceeded this way: \[5^{2(k+1)} -1 = 24q\] \[5^{2k+2}-1=24q\] \[5^{2k}5^{2}=24q+1\] \[5^{2k}25=24q+1\] then i got stuck

OpenStudy (anonymous):

oh nevermind! i didn't see your last reply before i sent that! thank you :)

ganeshie8 (ganeshie8):

np :) let me knw if smthng is not clear